THE PERIMETER OF A BASKETBALL COURT IS 108 METERS AND THE LENGTH IS 6 METERS longer than TWICE THE WIDTH? wHAT ARE THE LENGTH AND WIDTH
Answer by rfer(16322)

P=2w+2l
P=2w+2(2w+6)=108
2w+4w+12=108
6w=96
w=16
2w+6=38